OpenStudy (anonymous):

You are to manufacture a rectangular box with 3 dimensions x, y and z, and volume v=64. Find the dimensions which minimize the surface area of this box.

OpenStudy (zarkon):

Looks like a Lagrange multiplier problem

OpenStudy (anonymous):

so what will i have to do?

OpenStudy (zarkon):

\[\nabla f=\lambda \nabla g\] where \[f(x,y,z)=2xy+2xz+2yz\] and \[g(x,y,z)=xyz-64\]

OpenStudy (anonymous):

zarkon is there anyway of contacting u other than openstudy

OpenStudy (zarkon):

why?

OpenStudy (anonymous):

i feel like u can usually help me with contest questions

OpenStudy (anonymous):

its cool if u dont want to be bothered

OpenStudy (zarkon):

ic...and I'm not on OS enough :)

OpenStudy (anonymous):

oh lol kk, @ what times do u usally come to openstudy....just wanna know when i can catch u

OpenStudy (zarkon):

I get e-mail notification when someone replies to a thread that I have replied to. so you could find one of my posts and reply to it. I will then get an email.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how did u get the f(x,y,z)=2xy+2xz+2yz ?

OpenStudy (zarkon):

if the dimensions are x,y,z then the surface area is just the sum of the areas of the 6 sides xy+xy+xz+xz+yz+yz=2xy+2xz+2yz

OpenStudy (anonymous):

ok i been trying to solve this but i can not this is what i have so far\[fx=2y+2z=\lambda(yz) fy= 2x+2z=\lambda(xz) fz= 2x+2y=\lambda(xy) \]

OpenStudy (zarkon):

ok so far

OpenStudy (zarkon):

do this... multiply the first one by x and the 2nd one by y to get \[2xy+2xz=\lambda xyz\] and \[2xy+2yz=\lambda xyz\] so \[2xy+2xz=2xy+2yz\] \[\[2xz=2yz\] \[x=y\]

OpenStudy (anonymous):

ok and we do the same for z right?

OpenStudy (anonymous):

i got z= y

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

so how will i get the value of any of them?

OpenStudy (anonymous):

v=64 should i 64/3?

OpenStudy (zarkon):

x=y=z so x*y*z=64 turns into \[x^3=64\] x=4

OpenStudy (anonymous):

ooo k

OpenStudy (anonymous):

thanx a lot of your help

OpenStudy (zarkon):

np